# SOUNDCOMPASS: NAVIGATING TARGET SOUND EXTRACTION WITH EFFECTIVE DIRECTIONAL CLUE INTEGRATION IN COMPLEX ACOUSTIC SCENES

### v1

@@ 1 Introduction
To address these limitations, we propose SoundCompass, a framework for effective directional clue integration. The key contributions of the proposed framework are as follows:
SPIN input feature To build a general input feature capturing full spatial information, we propose a Spectral Pairwise INteraction (SPIN) module that captures all pairwise interactions between sinusoidal components on complex spectrogram of multichannel signals. SPIN features are then fused with a DoA clue in overlapping frequency subbands to capture frequency-dependent spatial cues, extending the advantages of prior band-split approaches [19, 20, 21, 22].
SH embedding for DoA clues We employ spherical harmonics (SH) as the DoA clue embedding, which provides a continuous angular representation across the 2D sphere. The SH embedding enables the model to handle any DoA value without discretization.
Iterative refinement with temporal clue Furthermore, we adopt an iterative refinement strategy inspired by the chain-of-inference (CoI) paradigm [23, 24, 25], where the estimated temporal activation is recursively fused with the DoA clue, enabling the model to improve separation quality under challenging multi-source conditions progressively.
@@ 2.1 Model Architecture
Our proposed model is based on DeepASA [25] backbone, which achieved state-of-the-art performance in universal source separation (USS) and sound event localization and detection (SELD). The architecture sequentially applies multi-head self-attention and Mamba feedforward networks along spectral and temporal dimensions separately to capture object-level features from mixtures. The original backbone separates object features without any clue, and our focus here is how to effectively guide the object separation process using a DoA clue.
The overall architecture for achieving this objective is illustrated in Fig. 1(a), excluding the batch dimension. A multichannel mixture is first transformed into a complex spectrogram of shape $2M\times T\times F$ using the short-time Fourier transform (STFT), where $M$, $T$, and $F$ denote the number of microphones, time frames, and frequency bins, respectively. Instead of a fixed window, a learnable Gaussian window parameterized by adaptive mean and standard deviation is applied, allowing each frame to be spectralized from different temporal focus and spread. The spectrogram is then mapped into a feature space, increasing the channel dimension from $2M$ to $D$ through a Conv2D encoder (kernel size 3 and stride 1) that extracts spatial cues and local spectral–temporal patterns. The resulting feature is modulated by a fusion module to align extracted features with the directional clue, as described in section 2.2. Subsequently, fused features are processed by feature aggregation (FA) blocks that analyze spectral and temporal dependencies and separate features corresponding to the directional clue. Finally, two audio decoders reconstruct the multichannel direct sound and reverberation separately by reducing the channel dimension from $D$ back to $2M$ using Conv2D layers of kernel size 3 and stride 1, followed by an inverse STFT (iSTFT) to recover the target waveform.
@@ 2.2 Directional Clue Integration
Spectral Pairwise INteraction The fusion module takes the complex spectrogram and directional clues to generate the spatial feature mask that guides the DeepASA system to the target direction. To aid in extracting spatial details from the complex spectrogram, we present a SPIN module. Within this module, the cosine and sine components of individual microphone signals are multiplied in the complex spectrogram domain, yielding a channel dimension of $(2M)^{2}$. This multiplication enhances the recognition of inter-channel phase or time differences, as well as level differences when necessary. The sinusoidal products, confined within a range of $\pm 1$, ensure stable learning dynamics during training. Given that inter-channel relationships often vary by frequency, we adopt a band-split strategy [19, 20, 21, 22] to ensure that inter-channel features are developed and merged with directional clues in each frequency band. Specifically, we employ overlapping subbands based on the 12-TET Western musical scale [20, 22], using $K=31$ subbands, with narrower bandwidths at lower frequencies and wider bandwidths at higher frequencies. This approach promotes continuity across overlapping frequencies and minimizes information loss at subband boundaries.
Spherical harmonics embedding As an accurate and continuous representation of the DoA clue $(\theta,\phi)$, we employ spherical harmonics (SH) as embeddings, as depicted in Fig. 1(b). Unlike one-hot embeddings, SH embeddings allow for the representation of angles without the need for discretization. Additionally, in contrast to cyc-pos embeddings, which define azimuth and elevation angles separately, SH embeddings can describe the position on an $S^{2}$ sphere without coordinate separation, thus providing a consistent representation regardless of coordinate rotation. The complex spherical harmonics of order $n$ and degree $m$ are defined as
@@
where $P_{n}^{m}(\cdot)$ denotes associated Legendre functions. We use the 5-th order encoding by stacking the real and imaginary components of spherical harmonics, yielding an embedding vector of dimension $2(N+1)^{2}$ for $N=5$.
Fusion in subbands The encoded SH vector is fused with the output from the SPIN module in each subband. The clue encoding module is a FiLM [26] layer generating scale ($\gamma$) and shift ($\beta$) parameters for feature modulation, combined with a residual connection from the feature encoding module. This design enables fine-grained spatial conditioning without hand-crafted feature engineering. Each encoding and decoding block consists of a linear layer, adaptive layer normalization (AdaNorm) [27], and a parametric rectified linear unit (PReLU) activation.
Iterative refinement To enhance robustness, we incorporate a sound event detection (SED) decoder following the on/off decoder structure of DeepASA [25]. This module outputs a frame-wise binary sequence that indicates the presence of the target source at each time step. As illustrated in Fig. 2, this sequence is combined with the SH embedding to form a time-varying directional clue of shape $T^{\prime}\times 2(N+1)^{2}$, where $T^{\prime}$ denotes the sequence length of the SED decoder output. The clue is linearly interpolated to $T$ in the time dimension and recursively injected into subsequent TSE stages. This chain-of-inference (CoI) strategy enables the model to iteratively refine its extraction by aligning directional information with temporal dynamics, with no additional modules required. During training, the already trained first-stage model is kept fixed, and only the subsequent stage is fine-tuned from the first-stage model using oracle time-varying clues while fixing its encoder and fusion module. This fine-tuned stage then functions as the subsequent stage. At evaluation, the complete pipeline is used: the output of the first-stage model is combined with the SH embedding and fed into the subsequent stage, which relies on SED predictions from the previous stage to progressively refine the separation of the target source.
@@ 2.3 Loss Functions
A linear combination of signal-to-noise ratio (SNR) and scale-invariant signal-to-noise ratio (SI-SNR) [28] loss with a ratio of 9:1 was used for the direct/reverb audio decoder and the sum of outputs from the decoders. Additionally, binary cross-entropy (BCE) loss was employed for the SED decoder to estimate the temporal activation of the target source. All terms were summed with identical weights.
@@ 3.1 Datasets
The proposed architecture was trained and evaluated on the Auditory Scene Analysis V2 (ASA2) dataset [^1] https://huggingface.co/datasets/AnonymousAuthor1/ASA2_dataset [25], which contains 13 audio classes with 2–5 foreground sources and one background noise per mixture. However, to align with our direction-based TSE model, we regenerated the dataset with stationary sources using the gpuRIR library [^2] https://github.com/DavidDiazGuerra/gpuRIR [29], fixing each source at its initial position. Each mixture is 4 seconds long, sampled at 16 kHz, and recorded with a 4-channel tetrahedral microphone array of 4.2 cm radius centered in a cuboid room. The other configurations, including initial source positions and room reflections, follow those of the ASA2 dataset configuration. The resultant training, validation, and test sets comprise 50k, 2k, and 2k mixtures, respectively.
@@ 3.3 Analysis of Results
Ablation studies further highlight the contribution of each component. Removing pairwise interactions in the SPIN module (i.e., using only the raw $2M$ cosine and sine components without multiplication) causes a substantial degradation, highlighting the necessity of cross-channel correlation modeling. Replacing spherical harmonics (SH) with cyc-pos embeddings slightly degrades performance, while eliminating the band-split structure also reduces accuracy, underscoring the importance of frequency-dependent spatial cues. Incorporating an SED decoder shows modest improvements, while iterative refinement further boosts performance, demonstrating the advantage of progressively refining activations at the cost of additional parameters.
@@ 4 Conclusion
We proposed SoundCompass, a DoA-based TSE framework that integrates spherical harmonics encoding with spectral pairwise interaction for efficient spatial conditioning. Through overlapping band-split modulation and sound event activation estimation, the model effectively captures both frequency-dependent and time-varying spatial cues with low complexity. Furthermore, the iterative refinement strategy highlights the advantage of coupling DoA clues with temporal dynamics, enabling a more robust extraction in diverse conditions. These results suggest promising directions toward more flexible source manipulation, such as handling dynamically moving sources by jointly estimating their time-varying DoA and activity.

### v2

@@ 1 Introduction
To address these limitations, we propose SoundCompass, a framework for more effective directional clue integration. The key contributions of the proposed framework are as follows:
SPIN input feature. To build a general input feature capturing rich spatial information, we propose a Spectral Pairwise INteraction (SPIN) module that captures all pairwise interactions between sinusoidal components on complex spectrogram of multichannel signals. SPIN features are then fused with a DoA clue in overlapping frequency subbands to capture frequency-dependent spatial cues, extending the advantages of prior band-split approaches [12, 23, 26, 22].
SH embedding for DoA clues. We employ spherical harmonics (SH) as the DoA clue embedding, which provides a continuous angular representation across the 2D sphere. The SH embedding enables the model to handle any DoA value without discretization.
Iterative refinement with temporal clue. Furthermore, we adopt an iterative refinement strategy inspired by the chain-of-inference (CoI) paradigm [24, 8, 10], where the estimated temporal activation is recursively fused with the DoA clue, enabling the model to improve separation quality under challenging multi-source conditions progressively.
@@ 2.1 Model Architecture
Our proposed model is based on DeepASA [10] backbone, which achieved state-of-the-art performance in universal source separation (USS) and sound event localization and detection (SELD) on their benchmarks. The architecture sequentially applies multi-head self-attention and Mamba feedforward networks along spectral and temporal dimensions separately to capture object-level features from mixtures. The original backbone separates object features without any clue, and our focus here is how to effectively guide the object separation process using a DoA clue.
The overall architecture for achieving this objective is illustrated in Fig. 1(a), excluding the batch dimension. A multichannel mixture is first transformed into a complex spectrogram of shape $2M\times T\times F$ using the short-time Fourier transform (STFT), where $M$, $T$, and $F$ denote the number of microphones, time frames, and frequency bins, respectively. Instead of a fixed window, a learnable Gaussian window parameterized by adaptive mean and standard deviation is applied, allowing each frame to be spectralized from different temporal focus and spread. The spectrogram is then mapped into a feature space, increasing the channel dimension from $2M$ to $D$ through a 2D convolutional encoder (kernel size 3 and stride 1) that extracts spatial cues and local spectral–temporal patterns. The resulting feature is modulated by a fusion module to align extracted features with the directional clue, as described in section 2.2. Subsequently, fused features are processed by feature aggregation (FA) blocks that analyze spectral and temporal dependencies and separate features corresponding to the directional clue. Finally, two audio decoders reconstruct the multichannel direct sound and reverberation separately by reducing the channel dimension from $D$ back to $2M$ using 2D convolutional layers of kernel size 3 and stride 1, followed by an inverse STFT (iSTFT) to recover the target waveform.
@@ 2.2 Directional Clue Integration
Spectral Pairwise INteraction (SPIN). The fusion module takes the complex spectrogram and directional clues to generate the spatial feature mask that guides the DeepASA system to the target direction. To aid in extracting spatial details from the complex spectrogram, we present a Spectral Pairwise INteraction (SPIN) module. Within this module, the cosine and sine components derived from the phase of each multichannel complex STFT are multiplied pairwise across channels, yielding a channel dimension of $(2M)^{2}$. This multiplication enhances the recognition of inter-channel phase or time differences, as well as level differences when necessary. The sinusoidal products, confined within a range of $\pm 1$, ensure stable learning dynamics during training. Given that inter-channel relationships often vary by frequency, we adopt a band-split strategy [12, 23, 26, 22] to ensure that inter-channel features are developed and merged with directional clues in each frequency band. Specifically, we employ overlapping subbands based on the 12-TET Western musical scale [23, 22], using $K=31$ subbands, with narrower bandwidths at lower frequencies and wider bandwidths at higher frequencies. This approach promotes continuity across overlapping frequencies and minimizes information loss at subband boundaries.
Spherical harmonics (SH) embedding. As an accurate and continuous representation of the DoA clue $(\theta,\phi)$, we employ spherical harmonics (SH) as embeddings, as depicted in Fig. 1(b). Unlike one-hot embeddings, SH embeddings allow for the representation of angles without the need for discretization. Additionally, in contrast to cyc-pos embeddings, which define azimuth and elevation angles separately, SH embeddings can describe the position on an $S^{2}$ sphere without coordinate separation, thus providing a consistent representation regardless of coordinate rotation. The complex spherical harmonics of order $n$ and degree $m$ are defined as
@@
where $P_{n}^{m}(\cdot)$ denotes associated Legendre functions. We used up to the 5-th order encoding by stacking the real and imaginary components of spherical harmonics, yielding an embedding vector of dimension $2(N+1)^{2}$ for $N=5$.
Fusion in subbands. The encoded SH vector is fused with the output from the SPIN module in each subband. The clue encoding module is a FiLM [19] layer generating scale ($\gamma$) and shift ($\beta$) parameters for feature modulation, combined with a residual connection from the feature encoding module. This design enables fine-grained spatial conditioning without hand-crafted feature engineering. Each encoding and decoding block consists of a linear layer, adaptive layer normalization (AdaNorm) [25], and a parametric rectified linear unit (PReLU) activation.
Iterative refinement. To enhance robustness, we incorporate a sound event detection (SED) decoder following the on/off decoder structure of DeepASA [10]. This module outputs a frame-wise binary sequence that indicates the presence of the target source at each time step. As illustrated in Fig. 2, this sequence is combined with the SH embedding to form a time-varying directional clue of shape $T^{\prime}\times 2(N+1)^{2}$, where $T^{\prime}$ denotes the sequence length of the SED decoder output. The clue is then linearly interpolated to $T$ in the time dimension and recursively injected into subsequent TSE stages. This chain-of-inference (CoI) strategy enables the model to iteratively refine its extraction by aligning directional information with temporal dynamics. During training, the already trained first-stage model is kept fixed, and only the subsequent stage is fine-tuned from the first-stage model using oracle time-varying clues while fixing its encoder and fusion module. This fine-tuned stage then functions as the subsequent stage. At evaluation, the complete pipeline is used: the output of the first-stage model is combined with the SH embedding and fed into the subsequent stage, which relies on SED predictions from the previous stage to progressively refine the separation of the target source.
@@ 2.3 Loss Functions
A linear combination of signal-to-noise ratio (SNR) and scale-invariant signal-to-noise ratio (SI-SNR) [9] loss with a ratio of 9:1 was used for the direct/reverb audio decoder and the sum of outputs from the decoders. Additionally, binary cross-entropy loss was employed for the SED decoder to estimate the temporal activation of the target source. All terms were summed with identical weights.
@@ 3.1 Datasets
The proposed architecture was trained and evaluated on the Auditory Scene Analysis V2 (ASA2) dataset [^1] https://huggingface.co/datasets/donghoney22/ASA2_dataset [10], which contains 13 audio classes with 2–5 foreground sources and one background noise per mixture. However, to align with our direction-based TSE model, we regenerated the dataset with stationary sources using the gpuRIR library [^2] https://github.com/DavidDiazGuerra/gpuRIR [2], fixing each source at its initial position. Each mixture is 4 seconds long, sampled at 16 kHz, and recorded with a 4-channel tetrahedral microphone array of 4.2 cm radius centered in a cuboid room. The other configurations, including initial source positions and room reflections, follow those of the ASA2 dataset configuration. The resultant training, validation, and test sets comprise 50k, 2k, and 2k mixtures, respectively.
@@ 3.3 Analysis of Results
Ablation studies further highlight the contribution of each component. Removing pairwise interactions in the SPIN module (i.e., using only the raw $2M$ cosine and sine components without multiplication) causes a substantial degradation, highlighting the importance of cross-channel correlation modeling. Replacing spherical harmonics (SH) with cyc-pos embeddings slightly degrades performance, while eliminating the band-split structure also reduces accuracy, underscoring the importance of frequency-dependent spatial cues. Incorporating an SED decoder shows modest improvements, while iterative refinement further boosts performance, demonstrating the advantage of progressively refining activations at the cost of additional parameters.
@@ 4 Conclusion
We proposed SoundCompass, a DoA-based TSE framework that integrates spherical harmonics embedding with spectral pairwise interaction for efficient spatial conditioning. Through overlapping band-split modulation and sound event activation estimation, the model effectively captures both frequency-dependent and time-varying spatial cues with low complexity. Furthermore, the iterative refinement strategy highlights the advantage of coupling DoA clues with temporal dynamics, enabling a more robust extraction in diverse conditions. These results suggest promising directions toward more flexible source manipulation, such as handling dynamically moving sources by jointly estimating their time-varying DoA and activity.